

********************************************************************
UARS WEEKLY STATUS REPORT - 12 March 2004
********************************************************************
A real good friend and giant in our field and a great person
has passed away. A sad day for our community.
Jim Holton, professor of atmospheric sciences at the University
of Washington/Seattle, passed away Wednesday (March 3) around
8:30 pm PST. Jim suffered a major heart attack while running
last week and apparently never regained consciousness.
Jim Holton was a Principal Investigator for UARS.
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
51% of each day's data amounting to about 510 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1994351. MEPS data files between
1994352 and 1994365 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. These is additional data for the
October and November storms of 2003 located at
http://amie.space.swri.edu/UARS_images_from_Rudy_(Feb._2004).html.
For those users of SDDAS, update your UARS Meta data and they will appear
with full functionality.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the Mini-Operational Readiness
Review meeting held on 9 March in regard to the use of
three TDRSS contacts on each UARS orbit. Testing by
FOT engineers has been successful. Routine use of three
TDRSS contacts per orbit is expected to begin on
17 March 2004. This change will increase the amount of
data acquired by UARS instruments, and reduce reliance
on the tape recorder for instances of critical data collection.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 28 February 2004 (Orbit 68202, GMT Day 059) through
05 March 2004 (Orbit 68307, GMT Day 065).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 061/1956 68244 ZEPS/NEPS TO DAY ONLY
HALOE 065/1516 68302 POWER OFF
The beta angle ranged between the angles of 26.8 degrees to 50.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 60 degrees on 11 March 2004 (DOY 071).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 2.23 to 1.90 degrees C. Additional
battery performance data for each battery on 27 February and 05 March
2004 are provided in Appendix D.
The clock error ranged from 10.1 to -09.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 01
HALOE = 01
PEM = 01
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continued testing the new version of the CMS code this week.
The testing is scheduled to be completed by 05 March 2004. A
tentative date of 09 March 2004 has been set for a mini-ORR to review
the results of testing and the plans to utilize the new 3-TDRS
functionality.
MEETINGS
The FOT attended the UARS Monthly Status Review on 03 March 2004 to
review the significant events from the month of February.
The FOT also attended the biweekly Eclipse DRB to discuss open DRs and
the status of the current build.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 March 2004 (DOY 089). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 March 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68103 052/1031 04-062 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68105 052/1353 04-061 MACS YAW -FWD TO REV
68121 053/1501 04-061 MPS VT LEVEL 5 (NORMAL)
68121 053/1545 04-061 MPS BATSOC
68123 053/1835 04-063 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
68134 054/1220 04-064 SOLSTICE FULLTIME OPS
68136 054/1542 04-065 SUSIM POWER ON (BEGIN)
68151 055/1539 04-066 SUSIM POWER ON (COMPLETE)
68165 056/1356 04-067 HALOE POWER ON
68165 056/1356 04-068 HALOE POWER OFF
68166 056/1453 04-067 HALOE POWER ON
68168 056/1829 04-069 HALOE POWER ON (PART 2)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 March 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 March 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68103 052/1034 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
68123 053/1835 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 March 2004
27 February 2004 - GMT Day 058 Beta = 25.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +16.8/-072.8 +5.46/+4.50 +28.8/-14.4 25.6 1.03 20.5
3 +11.2/-112.0 +4.50/+3.23 +28.4/-14.4 25.4 1.02 19.9
05 March 2004 - GMT Day 065 Beta = 50.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +16.8/-100.8 +5.14/+2.91 +25.2/-14.8 25.8 1.09 17.2
3 +11.2/-095.2 +3.87/+1.97 +24.8/-15.2 25.6 1.08 17.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None This Report Period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 March 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-011 FOT 68230 060/2042 30832 Generic Late Acquisition
04-014 FOT 68293 060/0126 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-011 FOT See List Feb 04 30832 Generic Late Acquisition
04-013 FOT 68165 056/1311 35212 Multiple Data Dropouts
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-011 FOT See List Feb 04 30832 Generic Late Acquisition
04-014 FOT See List Mar 04 30832 Generic Late Acquisition